9. A computer programmer has a 40% chance of finding a bug in any given program. What is the (1 point, probability that she finds her first bug in the sixth program she examines? 00.97 00.95 00.17 00.03 10. Donnell is doing market research for a major brand by surveying customers at a local mall. ( point) The probability of a customer agreeing to take the survey is 22%, what is the probability that the fourth customer is the first one to take the survey 010% 013% О 21% 63%
Answer #1

Answer :

9. Probability that first bug is found in 6th program = probability that first five programs are bug free × probability that 6th program contains bug

Probability of finding bug = 0.40

Probability of not finding bug = 0.60

Required probability = (0.60)^5 * 0.40 = 0.031104

So answer : d) 0.03

10. Probability of customer agreeing to take survey := 0.22

Probability of customer not agreeing to take survey = 0.78

Probability that 4th customer is first one to take survey = probability that fist three customers do not agree to take survey * probability that 4th customer take survey = 0.78*0.78*0.78*0.22 = 0.1044 = 10.44%

Answer : a) 10%

Please upvote